OpenStudy (anonymous):

Solve for X. X/4 = 2 1/4 *

OpenStudy (thadds2003):

What grade are you in?

OpenStudy (michele_laino):

we can rewrite he right side as below: \[2\frac{1}{4} = 2 + \frac{1}{4} = \frac{{\left( {2 \times 4} \right) + 1}}{4} = ...?\]

OpenStudy (michele_laino):

the*

OpenStudy (michele_laino):

please complete the computation

OpenStudy (michele_laino):

hint: \[2\frac{1}{4} = 2 + \frac{1}{4} = \frac{{\left( {2 \times 4} \right) + 1}}{4} = \frac{{8 + 1}}{4} = \frac{9}{4}\] so the original equation can be rewritten as below: \[\frac{x}{4} = \frac{9}{4}\] next I multiply both sides by 4: \[\frac{x}{4} \cdot 4 = \frac{9}{4} \cdot 4\] please simplify

OpenStudy (anonymous):

@Michele_Laino im back

OpenStudy (michele_laino):

so, what is the final result?

OpenStudy (anonymous):